OpenStudy (anonymous):

Find the area for each circle with a radius of 32 yd. A. 3200 yd^2 B. 3215.4 yd^2 C. 3222 yd^2 D. 3229.3 yd^2

OpenStudy (agent0smith):

\[\Large Area = \pi \times r^2\] r=radius, which is 32yd.

OpenStudy (anonymous):

lol @Luis_Rivera yep agreed :P

OpenStudy (agent0smith):

lol yeah i wondered about that too

OpenStudy (anonymous):

dude who knows its a virtual highschool website

OpenStudy (anonymous):

\[A = \pi \times r^2\] \[A = \pi \times 32^2\] \[A = \pi \times 1024\] \[A = 1024\pi\]

OpenStudy (anonymous):

A = 1024pi yds^2

OpenStudy (anonymous):

A = 1024 * 3.14 A = 3215.36 \[A \approx 3215.4\]